# Information-Theoretic Reward Decomposition for Generalizable RLHF

### v1

@@ Abstract
A generalizable reward model is crucial in Reinforcement Learning from Human Feedback (RLHF) as it enables correctly evaluating unseen prompt-response pairs. However, existing reward models lack this ability, as they are typically trained by increasing the reward gap between chosen and rejected responses, while overlooking the prompts that the responses are conditioned on. Consequently, when the trained reward model is evaluated on prompt-response pairs that lie outside the data distribution, neglecting the effect of prompts may result in poor generalization of the reward model. To address this issue, we decompose the reward value into two independent components: prompt-free reward and prompt-related reward. Prompt-free reward represents the evaluation that is determined only by responses, while the prompt-related reward reflects the reward that derives from both the prompt and the response. We extract these two components from an information-theoretic perspective, which requires no extra models. Subsequently, we propose a new reward learning algorithm by prioritizing data samples based on their prompt-free reward values. Through toy examples, we demonstrate that the extracted prompt-free and prompt-related rewards effectively characterize two parts of the reward model. Further, standard evaluations show that our method improves both the alignment performance and the generalization capability of the reward model.
Machine Learning, ICML
Figure 1: Left: reward gaps calculated with corresponding prompt and randomly sampled prompts using QRM-Llama3-8B 11footnotemark: 100footnotetext: xx on two different datasets that were used for training. When calculating with other prompts, the curves show the mean and the std. Right: illustrative failure case where the reward gap overly depends on the responses. Solid lines represent corresponding prompt-response pairs (used in training), while dashed lines represent non-corresponding pairs (unseen during training). Since the reward gap overly depends on the responses, it generalizes poorly to novel prompt-response pairs constructed even with seen but non-corresponding prompts.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is an effective approach for Large Language Models (LLMs) alignment (Christiano et al., 2017; Bai et al., 2022). Within a wide range of RLHF methods, reward learning plays a pivotal role. These methods typically first train a reward model on a static dataset and then leverage it to do Reinforcement Learning (RL) (Ouyang et al., 2022; Dong et al., 2024). Compared with methods that are free of using reward models (Rafailov et al., 2024b; Zhao et al., 2023), the advantage of such methods is their capacity to leverage the generalization capability of the reward model to evaluate outside-of-distribution prompt-response pairs. These prompt-response pairs with generated rewards can be used to further improve the LLM’s performance (Stiennon et al., 2020; Ziegler et al., 2019).
@@
Clearly, learning a generalizable reward model is central to this scenario. However, we found that standard reward training does not guarantee sufficient generalization capability. In reward model training, the primary goal is typically better distinguishing between chosen and rejected responses. To achieve this, the reward model does not necessarily require consideration of the corresponding prompt. Taking reward learning based on Bradley-Terry (BT) model as an example. Since the potential response space is vastly larger than the dataset size, different data samples typically contain distinct response pairs. As long as the reward gap within each pair increases, the training loss will decrease effectively. This can occur even if the reward model only takes the responses as the input. For the reward model that separately calculates rewards for chosen and rejected prompt-response pairs, the situation becomes similar as the inputs to the reward model only differ starting from the response.
@@
Perhaps surprisingly, such a phenomenon indeed appears in current reward models, even those that achieve SOTA performance on benchmarks. As shown in Fig. 1 (left), after replacing the corresponding prompt with other prompts, the reward gaps still center around their original values. This issue, where responses dominate the reward gap, does not affect training but leads to catastrophic results when evaluating novel prompt-response pairs. The illustrative example in Fig. 1 (right) shows this. When considering each prompt-response pair separately within the training dataset, the reward model appears to perform perfectly. However, when querying preferences after replacing the original prompt with other even in-distribution prompts (which are also meaningful queries), the reward model can yield significantly inaccurate or even reversed preferences. This generalization issue will become more pronounced when dealing with unseen prompt-response pairs encountered during evaluation. All of these highlight the need to distinguish two components of the reward value: one part is the value determined solely by the response, and the other is the value that can only be derived by simultaneously considering both the prompt and the response. We refer to the former as prompt-free reward and the latter as prompt-related reward.
@@
To address this, we propose a novel method of decomposition to extract these two components from an information-theoretic perspective, without requiring extra models. After that, we use the extracted prompt-free reward to guide the reward learning process, prioritizing training samples based on their prompt-free reward gap values. We verify our method through several toy examples and standard evaluations based on commonly used datasets and base models. In toy examples, the extracted prompt-free and prompt-related reward gaps characterize the reward model’s preference bias about response-only features and the generalizable preference information, respectively. Moreover, in standard experiments with common datasets, the reward model trained with our method outperforms strong baselines. These experiments show that considering both prompt-free and prompt-related rewards during training enhances the alignment performance and generalization capabilities of the reward model. footnotetext: 1https://huggingface.co/nicolinho/QRM-Llama3-8B.
@@ 2 Preliminaries
Standard preference learning assumes the existence of the preference oracle which determines $\mathbb{P}(y_{w}\succ y_{l}|x)$, where the probability of response $y_{w}$ is more preferred than $y_{l}$ conditioned on the prompt $x$. Given a preference dataset $D=\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$ where the prompt-response tuples $(x,y_{w},y_{l})$ are generated following $\mathbb{P}(y_{w}\succ y_{l}|x)$, our objective is to estimate the preference oracle from it.
There exist several methods to model the preference oracle (Luce, 1959), among which the Bradley-Terry (BT) model (Bradley & Terry, 1952) is the most widely used. The BT model further assumes that the preference oracle can be represented as a reward model $r:(x,y)\to\mathbb{R}$ that satisfies:
@@
Under this assumption, standard methods leverage a parameterized reward model $r_{\theta}$ to perform maximum likelihood estimation on the preference dataset (Ouyang et al., 2022; Stiennon et al., 2020) via the following objective:
@@
In this work, we only consider the preference learning based on the BT model, owing to its widespread use and strong performance in the field of RLHF. The parameterized reward model $r_{\theta}$ can be implemented in various ways (Touvron et al., 2023; Dong et al., 2024; Wang et al., 2024; Ivison et al., 2023). Among these, the most commonly used one is the ‘Sequence Classifier’ (Lambert et al., 2024; Ouyang et al., 2022), which builds the reward model on top of an LLM backbone. By projecting the feature representation of the prompt-response sequence to a scalar value using a simple network (e.g., single linear layer), such a reward model effectively leverages the prior knowledge embedded in the LLM backbone to evaluate prompt-response pairs. In this work, we focus on reward models within this type.
@@ 3 Methods
In this section, we first present a specific form of decomposition that divides the reward value into a free reward and a related reward. We then formalize the problem of such decomposition from a given reward model and propose a solution from an MI perspective. We also demonstrate that the solution can be effectively obtained without extra models.
@@ 3.1 Prompt-free Reward & Prompt-related Reward
We decompose the reward value $r_{\theta}(x,y)$ into two separate parts: the prompt-free reward and the prompt-related reward. The prompt-free reward is only determined by the response and can be regarded as the overall evaluation of the response. Specifically, the prompt-free reward ($r_{2}$) satisfies Eq. (3). Once the response is given, the prompt-free reward remains unaffected by any specific prompt, as
@@
Examining the dataset holistically, $r_{2}(x,y)$ demonstrates clear randomness, with the data distribution influencing its value. Specifically, for a given $(x,y)$ pair, if the prompts related to $y$ are diverse, there is little connection between $r_{2}(x,y)$ and $r_{\theta}(x,y)$, since $r_{2}$ remains unchanged across different prompts, whereas $r_{\theta}$ changes accordingly. Conversely, if the related prompts exhibit little variety, $r_{2}(x,y)$ can be closely related to $r_{\theta}(x,y)$, as the variation in $r_{\theta}(x,y)$ is also limited. This motivates us to define $r_{2}$ by considering the randomness in the data distribution, treating $r_{2}$ as the joint product of $r_{\theta}$ and the data distribution.
To obtain the overall evaluation of $r_{\theta}$ on $y$, one may consider simple transformations like directly taking the expectation of $r_{\theta}(x,y)$ over the prompt distribution, i.e. $r_{2}(x,y)=\mathbb{E}_{x^{\prime}\sim P(X|Y=y)}[r_{\theta}(x^{\prime},y)]$. Although $r_{2}(x,y)$ defined in this way is independent of $x$, such a definition can be inconsistent with ‘prompt-free reward’ from the preference perspective. To illustrate this, consider two responses, $y_{1}$ and $y_{2}$. If $r_{\theta}(x,y_{1})$ is significantly larger than $r_{\theta}(x,y_{2})$ under a single prompt, then even if $r_{\theta}(x,y_{2})$ is slightly larger than $r_{\theta}(x,y_{1})$ under all other prompts, $r_{2}(x,y_{1})$ will be much larger than $r_{2}(x,y_{2})$. Such a prompt-free reward strongly favors $y_{1}$, despite $r_{\theta}$ overall preferring $y_{2}$. A detailed numerical example is given in Appendix B. This indicates that a carefully designed objective is required to induce the prompt-free reward.
3.2 Solving Prompt-free Reward via a MI Objective
Although directly solving $r_{2}(x,y)$ by marginalizing $r_{\theta}(x,y)$ is challenging, we can extract $r_{2}(x,y)$ from $r_{\theta}(x,y)$ via a preference perspective. For simplicity, we have notations
$\Delta r(x,y_{1},y_{2})$ $\coloneqq r(x,y_{1})-r(x,y_{2})$ (5)
$\sigma(\Delta r(x,y_{1},y_{2}))$ $\coloneqq 1/\big{(}1+\exp(-\Delta r(x,y_{1},y_{2}))\big{)}.$ (6)
@@ 3.2 Solving Prompt-free Reward via a MI Objective
Meanwhile, since both $r_{1}$ and $r_{2}$ must be applicable to all $(x,y_{1},y_{2})$, it is essential to take into account the entire dataset when considered from the preference perspective. The preferences of $r_{1}$, $r_{2}$, and $r_{\theta}$ over the entire dataset are inherently random and can be characterized by the following random variables:
$\tilde{W}$ $=\text{Ber}(\mathbb{E}_{x\sim P(X|Y_{1},Y_{2})}[\mathbb{P}_{r_{\theta}}(Y_{1}\succ Y_{2}|x)])$
$W$ $=\text{Ber}\big{(}\mathbb{P}_{r_{\theta}}(Y_{1}\succ Y_{2}|X)\big{)}=\text{Ber}\big{(}\sigma(\Delta r_{\theta}(X,Y_{1},Y_{2}))\big{)},$
where ‘Ber’ stands for Bernoulli random variable. Among them, $W$, $Z$, and $\tilde{Z}$ are the random preferences of their corresponding rewards. $\tilde{W}$ is the prompt-free random preference of $r_{\theta}$. For all these random variables, the randomness comes from the prompt, responses, and the Bernoulli distribution. We also provide a detailed interpretation of these random variables in Appendix D.1.
@@
$\begin{cases}\text{MI}(Z~{}\|~{}\tilde{W})=0,\hfill\\ \text{MI}(\tilde{Z}~{}\|~{}\tilde{W})=\text{MI}(\tilde{Z}~{}\|~{}W).\hfill\end{cases}$ (9)
@@
Figure 2: Sub-figure (a) illustrates the information in $W$ and $\tilde{W}$. Sub-figures (b) and (c) characterize the desired and undesired pairs of $Z$ and $\tilde{Z}$, respectively.
Although these conditions effectively constrain the information in $Z$ and $\tilde{Z}$, a trivial solution exists. Assume $\Delta r_{\theta}(x,y)$ is bounded for any prompt $x$ and response $y$. Consider an ill-formed $\Delta r_{1}$ such that for any prompt $x$ and response $y$, $\Delta r_{1}(x,y)=+\infty$. Then by the definition that $\Delta r_{\theta}(x,y)=\Delta r_{1}(x,y)+\Delta r_{2}(x,y)$, $\Delta r_{2}(x,y)=-\infty$ for any prompt $x$ and response $y$. In this case, $Z$ will always be $1$, and $\tilde{Z}$ will always be 00. This makes $\text{MI}(Z\|\tilde{W})=\text{MI}(\tilde{Z}\|\tilde{W})=\text{MI}(\tilde{Z}\|W)=0$ since $Z$ and $\tilde{Z}$ are both constants. Although these ill-formed $r_{1}$ and $r_{2}$ satisfy the conditions in Eq. (9), they provide no information of prompt-related or prompt-free preferences. This ill-formed example indicates that the MI condition in Eq. (9) is insufficient to constrain the information in $Z$ or $\tilde{Z}$ themselves. When $Z$ and $\tilde{Z}$ become constants, the information within them is lost, along with their MI with other random variables. To avoid this, we optimize the following constrained objective to derive $r_{1}$ and $r_{2}$ instead:
@@
In addition to the previous conditions, we maximize the information within $Z$ to ensure a diverse prompt-related preference, considering that $Z$ represents the random prompt-related preference across the entire dataset distribution. This constrained optimization problem is difficult to solve directly, and direct solutions may require additional parameterized reward models. However, we provide an efficient algorithm that obtains $\Delta r_{1}(x,y_{1},y_{2})$ and $\Delta r_{2}(x,y_{1},y_{2})$ without additional reward model. To see this, we first note that the second constraint is satisfied with a specific structure of $r_{2}$. Formally, the following theorem holds:
@@
When the value of $r_{2}$ depends only on the response, i.e. $r_{2}(x,y)=r_{2}(y)$, MI$(\tilde{Z}\|\tilde{W})$ $=$ MI$(\tilde{Z}\|W)$.
@@
The proof is given in Appendix D.2. Intuitively, this requires that the prompt-free part of $r_{1}^{*}$ could not distinguish between the responses within any given response pair. This aligns with our desiderata, as $r_{1}^{*}$ should only reflect the evaluations considering both the prompt and response. Although neither of the theorems explicitly shows the connection with the preference of $r_{\theta}$, we note that the feasibility condition requires $\Delta r_{1}^{*}(x,y_{1},y_{2})+\Delta r_{2}^{*}(x,y_{1},y_{2})$ to equal $\Delta r_{\theta}^{*}(x,y_{1},y_{2})$, which in turn avoids trivial solution. Theorem 2 also provides an efficient way to obtain $\Delta r_{1}^{*}(x,y_{1},y_{2})$ and $\Delta r_{2}^{*}(x,y_{1},y_{2})$. After replacing $\Delta r_{1}^{*}(x,y_{1},y_{2})$ with $\Delta r_{\theta}(x,y_{1},y_{2})-\Delta r_{2}^{*}(x,y_{1},y_{2})$ in the optimal solution in Theorem 2, $\Delta r_{2}^{*}(x,y_{1},y_{2})$ satisfies that for any $(y_{1},y_{2})\sim P(Y_{1},Y_{2})$:
$\underset{x\sim P(X|Y_{1}=y_{1},Y_{2}=y_{2})}{\mathbb{E}}[\sigma(\Delta r_{\theta}(x,y_{1},y_{2})-\Delta r_{2}^{*}(y_{1},y_{2}))]=\frac{1}{2}.$ (11)
Note that we replace $\Delta r_{2}^{*}(x,y_{1},y_{2})$ with $\Delta r_{2}^{*}(y_{1},y_{2})$ due to Theorem 1. Such $\Delta r_{2}^{*}(y_{1},y_{2})$ represents a consistent shift across different prompts. By subtracting this shift from $\Delta r_{\theta}(x,y_{1},y_{2})$, we can remove the prompt-free preference in it. It is straightforward to prove that the expectation in Eq. (11) decreases monotonically with the increase of $\Delta r_{2}^{*}(y_{1},y_{2})$, as shown in Appendix D.2. Combined with the bounded assumption of $r_{\theta}$, $\Delta r_{2}^{*}(y_{1},y_{2})$ can be obtained via binary search in a finite interval. The searched $\Delta r_{2}^{*}(y_{1},y_{2})$ can be interpreted as the ‘weighted average’ of $\Delta r_{\theta}(x,y_{1},y_{2})$, based on $P(X|Y_{1}=y_{1},Y_{2}=y_{2})$. Due to limited space, we provide the pseudo-code of the binary search process in Appendix C.1 (Alg. 1).
@@
Here $P(x|y_{1},y_{2})$ is the abbreviation of $P(X=x|Y_{1}=y_{1},Y_{2}=y_{2})$, so as other probabilities. Given a $(y_{1},y_{2})$, we first sample from $P(X)$, the marginal distribution of prompts, and then re-weight the samples using $P(y_{1},y_{2}|x)$. Since this value represents the generation probability of the responses for a specific prompt, it can be efficiently computed in some cases or approximated in others. We list the sampling schemes for different cases in Appendix C.2.

### v2

@@ Abstract
A generalizable reward model is crucial in Reinforcement Learning from Human Feedback (RLHF) as it enables correctly evaluating unseen prompt-response pairs. However, existing reward models can lack this ability, as they are typically trained by increasing the reward gap between the chosen and rejected responses, while overlooking the prompts that the responses are conditioned on. Consequently, when the trained reward model is evaluated on prompt-response pairs that lie outside the data distribution, neglecting the effect of prompts may result in poor generalization of the reward model. To address this issue, we decompose the reward value into two independent components: prompt-free reward and prompt-related reward. Prompt-free reward represents the evaluation that is determined only by responses, while the prompt-related reward reflects the reward that derives from both the prompt and the response. We extract these two components from an information-theoretic perspective, which requires no extra models. Subsequently, we propose a new reward learning algorithm by prioritizing data samples based on their prompt-free reward values. Through toy examples, we demonstrate that the extracted prompt-free and prompt-related rewards effectively characterize the two parts of the reward value. Further, standard evaluations show that our method improves both the alignment performance and the generalization capability of the reward model.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is an effective approach for Large Language Models (LLMs) alignment [8, 5]. Within a wide range of RLHF methods, reward learning plays a pivotal role. These methods typically first train a reward model on a static dataset and then leverage it to do Reinforcement Learning (RL) [28, 11]. Compared with methods that are free of using reward models [30, 40], the advantage of such methods is their capacity to leverage the generalization capability of the reward model to evaluate outside-of-distribution prompt-response pairs. These prompt-response pairs with generated rewards can be used to further improve the LLM’s performance [36, 42].
@@
Clearly, learning a generalizable reward model is central to this scenario. However, we found that standard reward training does not guarantee sufficient generalization capability. In reward model training, the primary goal is typically better distinguishing between chosen and rejected responses. To achieve this, the reward model does not necessarily require consideration of the corresponding prompt. Taking reward learning based on Bradley-Terry (BT) model as an example. Since the potential response space is vastly larger than the dataset size, different data samples typically contain distinct response pairs. As long as the reward gap within each response pair increases, the training loss will decrease effectively. This can occur even if the reward model only considers the responses and totally ignores the prompts. In this case, the trained reward model loses its generalization capability over different prompts and may exhibit incorrect preference for novel prompt-response pairs.
@@
Perhaps surprisingly, such a phenomenon indeed appears in current reward models, even some that achieve SOTA performance on common benchmarks. As shown in Fig. 1 (left), after replacing the corresponding prompt with other prompts in the dataset, the reward gaps still center around their original values. This issue, where responses dominate the reward gap, does not affect training but leads to catastrophic results when evaluating novel prompt-response pairs. The illustrative example in Fig. 1 (right) shows this. When considering each prompt-response pair separately within the training dataset, its reward gap matches the ideal value. However, when querying preferences after replacing the original prompt with other prompts in the dataset (which are also meaningful queries), the reward model can yield inaccurate or even wrong preferences. This generalization issue will become more pronounced when dealing with unseen prompt-response pairs encountered during evaluation. All of these highlight the need to distinguish two components of the reward value: one part is the value determined solely by the response, and the other is the value that can only be derived by simultaneously considering both the prompt and the response. We refer to the former as prompt-free reward and the latter as prompt-related reward.
@@
To address this, we propose a novel method of decomposition to extract these two components from an information-theoretic perspective, without requiring extra models. After that, we use the extracted prompt-free reward to guide the reward learning process, prioritizing training samples based on their prompt-free reward gap values. We verify our method through several toy examples and standard evaluations based on commonly used datasets and base models. In toy examples, the extracted prompt-free reward gaps reflect the reward model’s preference bias about response-only features, while the prompt-related reward gaps capture its generalizable preference information. Moreover, in standard experiments with common datasets, the reward model trained with our method outperforms strong baselines. These experiments show that considering both prompt-free and prompt-related rewards during training enhances the alignment performance and generalization capabilities of the reward model.
Figure 1: Left: reward gaps calculated with corresponding prompt and randomly sampled prompts using QRM-Llama3-8B [^1] https://huggingface.co/nicolinho/QRM-Llama3-8B on two different datasets that were used for training. When calculating with other prompts, the curves show the mean and the std. Right: illustrative failure case where the reward gap overly depends on the responses. Solid lines represent corresponding prompt-response pairs (used in training), while dashed lines represent non-corresponding pairs (unseen during training). Since the reward gap overly depends on the responses, it generalizes poorly to novel prompt-response pairs constructed even with seen prompts.
@@ 2 Preliminaries
Standard preference learning assumes the existence of the preference oracle which determines $\mathbb{P}(y_{w}\succ y_{l}|x)$, the probability that response $y_{w}$ is more preferred than $y_{l}$ conditioned on the prompt $x$. Given a preference dataset $D=\{(x,y_{w},y_{l})_{i}\}_{i=1}^{N}$ where the prompt-response tuples $(x,y_{w},y_{l})$ are generated following $\mathbb{P}(y_{w}\succ y_{l}|x)$, our objective is to estimate the preference oracle from it.
There exist several methods to model the preference oracle [24], among which the Bradley-Terry (BT) model [6] is the most widely used. The BT model further assumes that the preference oracle can be represented as a reward model $r:(x,y)\to\mathbb{R}$ that satisfies:
@@
Under this assumption, standard methods leverage a parameterized reward model $r_{\theta}$ to perform maximum likelihood estimation on the preference dataset [28, 36] via the following objective:
@@
In this work, we consider the preference learning based on the BT model, owing to its widespread use and strong performance in the field of RLHF. The parameterized reward model $r_{\theta}$ can be implemented in various ways [37, 11, 38, 17]. Among these, the most commonly used one is the ‘Sequence Classifier’ [19, 28], which builds the reward model on top of an LLM backbone. By projecting the feature representation of the prompt-response sequence to a scalar value using a simple network (e.g., single linear layer), such a reward model effectively leverages the prior knowledge embedded in the LLM backbone to evaluate prompt-response pairs. In this work, we focus on reward models within this type. A discussion of other reward model structures is given in Appendix C.4.
@@ 3 Methods
In this section, we first present a specific form of decomposition that divides the reward value into a prompt-free reward and a prompt-related reward. We then formalize the problem of such decomposition and propose a solution from an MI perspective. Finally, we use the extracted prompt-free rewards to prioritize training samples and guide reward learning.
@@ 3.1 Prompt-free Reward & Prompt-related Reward
We decompose the reward value $r_{\theta}(x,y)$ into two separate parts: the prompt-free reward and the prompt-related reward. The prompt-free reward is only determined by the response and can be regarded as the overall evaluation of the response. Specifically, the prompt-free reward ($r_{2}$) satisfies Eq. (3). Once the response is given, the prompt-free reward remains unaffected by any specific prompt.
@@
Examining the dataset holistically, $r_{2}(x,y)$ demonstrates clear randomness, with the data distribution influencing its value. Specifically, for a given $(x,y)$ pair, if the prompts related to $y$ are diverse, there is little connection between $r_{2}(x,y)$ and $r_{\theta}(x,y)$, since $r_{2}$ remains unchanged across different prompts, whereas $r_{\theta}$ changes accordingly. Conversely, if the related prompts show little variety, $r_{2}(x,y)$ can be closely related to $r_{\theta}(x,y)$, since $r_{\theta}(x,y)$ also exhibits little change. This motivates us to define $r_{2}$ by considering the randomness in the data distribution, treating $r_{2}$ as the joint product of $r_{\theta}$ and the data distribution. We also note that although some simple definitions of $r_{2}$ may seem reasonable (e.g. $r_{2}(x,y)=\mathbb{E}_{x^{\prime}\sim P(X|Y=y)}[r_{\theta}(x^{\prime},y)]$), they can’t truly reflect $r_{\theta}$’s preference regarding the response alone (See Appendix B for more details).
3.2 Solving Prompt-free Reward via a Mutual Information (MI) Objective
In this section, we extract $r_{2}(x,y)$ from $r_{\theta}(x,y)$ via a preference perspective, with a carefully designed mutual information (MI) objective. A brief introduction to MI is provided in Appendix A. For simplicity, we have notations
$\Delta r(x,y_{1},y_{2})\coloneqq r(x,y_{1})-r(x,y_{2})\quad\sigma(\Delta r(x,y_{1},y_{2}))\coloneqq 1/\big(1+\exp(-\Delta r(x,y_{1},y_{2}))\big)$ (5)
@@ 3.2 Solving Prompt-free Reward via a Mutual Information (MI) Objective
Meanwhile, since both $r_{1}$ and $r_{2}$ must be applicable to all $(x,y_{1},y_{2})$, it is essential to take into account the entire dataset when decomposing from the preference perspective. The preference labels of $r_{1}$, $r_{2}$, and $r_{\theta}$ over the entire dataset are inherently random and can be characterized by the following random variables:
$Z\coloneqq\text{Ber}\big(\sigma(\Delta r_{1}(X,Y_{1},Y_{2})\big),$ $\tilde{Z}\coloneqq\text{Ber}\big(\sigma(\Delta r_{2}(X,Y_{1},Y_{2})\big)$ (7)
$\tilde{W}\coloneqq\text{Ber}(\mathbb{E}_{x\sim P(X|Y_{1},Y_{2})}[\sigma(\Delta r_{\theta}(x,Y_{1},Y_{2}))]),$ $W\coloneqq\text{Ber}\big(\sigma(\Delta r_{\theta}(X,Y_{1},Y_{2}))\big)$
where ‘Ber’ stands for Bernoulli random variable. Among them, $Z$, $\tilde{Z}$ and $W$ are the random preference labels of their corresponding rewards. $\tilde{W}$ is the random prompt-free preference label of $r_{\theta}$. For all these random variables, the randomness comes from the prompt, responses, and the Bernoulli distribution. We also provide a detailed interpretation of these random variables in Appendix D.1.
@@
$\text{MI}(Z~\|~\tilde{W})=0,\quad\text{MI}(\tilde{Z}~\|~\tilde{W})=\text{MI}(\tilde{Z}~\|~W).$ (8)
Figure 2: (a) Information in $W$ and $\tilde{W}$. (b) Desired information in $Z$ and $\tilde{Z}$. (c) Undesired information in $Z$ and $\tilde{Z}$.
@@
Although these conditions effectively constrain the information in $Z$ and $\tilde{Z}$, a trivial solution exists. Assume $\Delta r_{\theta}(x,y)$ is bounded for any prompt $x$ and response $y$. Consider an ill-formed $\Delta r_{1}$ such that for any prompt $x$ and response $y$, $\Delta r_{1}(x,y)=+\infty$. Then by the definition that $\Delta r_{\theta}(x,y)=\Delta r_{1}(x,y)+\Delta r_{2}(x,y)$, $\Delta r_{2}(x,y)=$-$\infty$ for any prompt $x$ and response $y$. In this case, $Z$ will always be $1$, and $\tilde{Z}$ will always be 0. This makes $\text{MI}(Z\|\tilde{W})=\text{MI}(\tilde{Z}\|\tilde{W})=\text{MI}(\tilde{Z}\|W)=0$ since $Z$ and $\tilde{Z}$ are both constants. Although these ill-formed $r_{1}$ and $r_{2}$ satisfy the conditions in Eq. (8), they provide no information of prompt-related or prompt-free preferences. This ill-formed example indicates that the MI condition in Eq. (8) is insufficient to induce a reasonable decomposition. When $Z$ and $\tilde{Z}$ become constants, the information within them is lost, along with their MI with other random variables. To avoid this, we optimize the following constrained objective to derive $r_{1}$ and $r_{2}$, which ensures sufficient information in $Z$ ($H$ stands for Shannon’s entropy):
@@
In addition to the previous conditions, we maximize the information within $Z$, considering that $Z$ represents the random prompt-related preference label across the entire dataset. This constrained optimization problem is difficult to solve directly, and direct solutions may require additional parameterized reward models. However, we provide an efficient algorithm that obtains $\Delta r_{1}(x,y_{1},y_{2})$ and $\Delta r_{2}(x,y_{1},y_{2})$ without additional reward model. To see this, we first note that the second constraint is satisfied with a specific structure of $r_{2}$. Formally, the following theorem holds:
@@
When the value of $r_{2}$ depends only on the response, i.e. $r_{2}(x,y)=r_{2}(y)$, $(\tilde{Z}\|\tilde{W}) == MI$$(\tilde{Z}\|W)$.
@@
The proof is given in Appendix D.3. Intuitively, this requires the prompt-free part of $r_{1}^{*}$ to equally prefer both responses in any given pair. This aligns with our desiderata, as $r_{1}^{*}$ should prefer some response only after being provided with the specific prompt. Although neither of the theorems explicitly shows the connection with $r_{\theta}$, we note that the feasibility condition requires $\Delta r_{1}^{*}(x,y_{1},y_{2})+\Delta r_{2}^{*}(x,y_{1},y_{2})$ to equal $\Delta r_{\theta}^{*}(x,y_{1},y_{2})$, which encodes $r_{\theta}$’s information. Theorem 2 also provides an efficient way to obtain $\Delta r_{1}^{*}(x,y_{1},y_{2})$ and $\Delta r_{2}^{*}(x,y_{1},y_{2})$. After replacing $\Delta r_{1}^{*}(x,y_{1},y_{2})$ with $\Delta r_{\theta}(x,y_{1},y_{2})-\Delta r_{2}^{*}(x,y_{1},y_{2})$ in the optimal solution in Theorem 2, $\Delta r_{2}^{*}(x,y_{1},y_{2})$ satisfies that for any $(y_{1},y_{2})\sim P(Y_{1},Y_{2})$:
$\Phi(y_{1},y_{2})\coloneqq\underset{x\sim P(X|Y_{1}=y_{1},Y_{2}=y_{2})}{\mathbb{E}}[\sigma(\Delta r_{\theta}(x,y_{1},y_{2})-\Delta r_{2}^{*}(y_{1},y_{2}))]=\frac{1}{2}.$ (10)
Note that we replace $\Delta r_{2}^{*}(x,y_{1},y_{2})$ with $\Delta r_{2}^{*}(y_{1},y_{2})$ due to Theorem 1. Because of this, $\Delta r_{2}^{*}(y_{1},y_{2})$ will not change while taking the conditional expectation. Moreover, it can be proved that $\Phi(y_{1},y_{2})$ in Eq. (10) decreases monotonically with the increase of $\Delta r_{2}^{*}(y_{1},y_{2})$, as shown in Appendix D.3. Combined with the bounded assumption of $r_{\theta}$, we have:
$\lim_{\Delta r_{2}^{*}(y_{1},y_{2})\to+\infty}\Phi(y_{1},y_{2})=0,\quad\lim_{\Delta r_{2}^{*}(y_{1},y_{2})\to-\infty}\Phi(y_{1},y_{2})=1$ (11)
With this, $\Delta r_{2}^{*}(y_{1},y_{2})$ can be obtained via binary search in a finite interval. The searched $\Delta r_{2}^{*}(y_{1},y_{2})$ can be interpreted as the ‘weighted average’ of $\Delta r_{\theta}(x,y_{1},y_{2})$, based on $P(X|Y_{1}=y_{1},Y_{2}=y_{2})$. Due to limited space, we provide the pseudo-code of the binary search process in Appendix C.1 (Alg. 1). Note that such an algorithm doesn’t require any extra parameterized reward models.
@@
Here $P(x|y_{1},y_{2})$ is the abbreviation of $P(X=x|Y_{1}=y_{1},Y_{2}=y_{2})$, so as other probabilities. Given a $(y_{1},y_{2})$, we first sample from $P(X)$, the marginal distribution of prompts, and then re-weight the samples using $P(y_{1},y_{2}|x)$. Its value represents the response generation probability and thus can be efficiently computed in some cases or approximated in others. Moreover, since $P(y_{1},y_{2}|x)$ is independent of the trained reward model, each dataset only needs to pre-sample prompts once and compute the probabilities once. We list the sampling schemes for different cases in Appendix C.2.